# Voost: A Unified and Scalable Diffusion Transformer for Bidirectional Virtual Try-On and Try-Off

### v1

@@ 1 Introduction
Figure 1: Attention map comparison— CatVTON [15] shows dispersed attention unrelated to the query point, indicating weak spatial grounding. In contrast, our model produces sharply localized maps that align well with the corresponding garment regions, demonstrating stronger relational understanding.
@@
Early approaches [27, 85, 21, 13] have adopted warping-based strategies, often combined with generative modeling, but have struggled to maintain garment fidelity and structural consistency.
Effectively adapting T2I diffusion models [66, 18] to virtual try-on requires strong image-based conditioning and accurate modeling of garment–person correspondence. To this end, various efforts have been made to improve spatial alignment and relational understanding. These include incorporating auxiliary objectives [41, 91], leveraging reference networks [92, 14, 84], or injecting conditioning signals through spatial concatenation [15]. Despite these efforts, existing methods often fail to establish precise garment–person correspondence. As illustrated in Fig. 1, the dispersed attention patterns reflect limited relational understanding, often resulting in lower visual quality and failure to faithfully reconstruct garment details.
@@
Remarkably, our unified diffusion transformer achieves state-of-the-art performance on both VTON and VTOFF benchmarks, surpassing strong task-specific baselines [41, 14, 91, 15, 78, 82] optimized individually for each task. We provide extensive qualitative and quantitative analyses that validate the effectiveness of our unified approach and its ability to model garment–person interaction across diverse scenarios.
@@ 2 Related work
Diffusion models have rapidly evolved to support image-conditioned generation beyond the text-to-image setting [66, 62]. Early methods such as Textual Inversion [19] and DreamBooth [67] enabled concept-specific synthesis via embedding injection or fine-tuning, but incurred high computational costs. Later approaches improved efficiency by leveraging pre-trained encoders like CLIP [63] and DINO [9, 59], as seen in IP-Adapter [86] and AnyDoor [12], which use image embeddings to guide generation. To improve spatial consistency, MasaCtrl [6] introduced mutual self-attention to preserve object identity. This was extended by reference-based methods [34, 11, 84, 14, 76] that inject structured visual features through dedicated networks. More recently, concatenation-based methods [15, 35, 75] have emerged as lightweight alternatives that directly feed conditioning signals without auxiliary modules. Additionally, image-conditioned diffusion has expanded beyond image-to-image generation to video [26, 34, 5, 71] and 3D synthesis [49, 45, 79], broadening its applicability.
@@
Early virtual try-on systems primarily relied on warping-based methods [80, 28, 20], where the main challenge was to deform and align the garment image to match the target person’s pose and body shape. With the emergence of Generative Adversarial Networks (GANs) [22, 38, 39], a series of GAN-based approaches [27, 83, 53, 36, 90, 1, 21, 85, 13] achieved more realistic image synthesis through adversarial learning. However, these models still struggled to preserve garment details and maintain structural consistency.
With the advancement of large-scale text-to-image diffusion models [64, 66, 68, 62, 18, 4], recent works [92, 2, 56, 24, 14, 84, 87, 69, 15, 91] have adopted pre-trained diffusion backbones as strong generative priors for virtual try-on. A key challenge remains: effectively guiding the model to learn the complex relationship between the garment and the target person. To address this, various techniques have been proposed to improve spatial correspondence and visual fidelity—such as auxiliary loss functions [41, 91], dual or reference networks [92, 14, 10], warping [56, 24, 90, 91], external image encoders [41, 14, 84], and spatial concatenation [15]. However, many of these approaches rely on weak garment–person coupling or introduce additional components that may degrade generation quality.
While most existing work focuses on virtual try-on, a few studies [78, 82] have explored virtual try-off, the inverse of virtual try-on, which aims to predict a clean garment image from a person wearing it. This task is inherently challenging due to occlusion, wrinkles, and deformation, and similarly requires strong relational modeling. We propose a unified DiT-based framework that jointly learns virtual try-on and try-off using a shared spatial conditioning layout. This design enables robust bidirectional garment–person modeling and remains scalable across diverse fashion imagery.
@@ 3.1 Preliminary
Diffusion models [32, 73, 66] generate images by denoising a sequence of latent representations corrupted over time. Given a latent $\mathbf{z}_{t}$, the network $\epsilon_{\theta}$ is trained to recover the clean signal $\mathbf{z}_{0}$ or the noise $\boldsymbol{\epsilon}$, conditioned on external input $\mathbf{y}$. In general, the denoising step follows:
@@
where ${\Phi}(\cdot)$ denotes an integration step, such as DDIM [72], DPM-Solver [52], or deterministic alternatives like flow matching [48]. In virtual try-on, $\mathbf{y}$ typically includes a garment image and a masked person image, optionally with pose or parsing maps [8, 25]. We adopt a unified DiT [60] that handles both try-on and try-off tasks. Details follow in Sec. 3.2.
@@
A key challenge in virtual try-on is aligning the garment structure with the target body while preserving visual fidelity. We assess this correspondence by analyzing transformer attention maps at specific spatial queries. As shown in Fig.1, CatVTON [15] exhibits dispersed attention across unrelated areas, indicating weak spatial grounding, which can reduce structural fidelity and impair detail preservation. In contrast, attention maps produced by our model are sharply localized and semantically consistent, suggesting stronger garment–target alignment. This motivates our bidirectional training strategy, where jointly learning try-on and try-off within a single model improves spatial alignment.
@@ 3.2 Try-On & Off via Unified Transformer
Let $\mathbf{X_{g}}\in\mathbb{R}^{H\times W\times 3}$ denote the standalone garment image, and $\mathbf{X_{p}}\in\mathbb{R}^{H\times W\times 3}$ the target person image. We construct the input by horizontally concatenating the two, forming a combined image $\mathbf{X}=[\mathbf{X_{g}}\mid\mathbf{X_{p}}]\in\mathbb{R}^{H\times 2W\times 3}$. Task-specific inpainting regions are defined by a binary mask $\mathbf{M}\in\{0,1\}^{H\times 2W}$, applied in the image space prior to encoding. For the try-on task, the mask is defined as $\mathbf{M}=[\mathbf{0}\mid\mathbf{M_{\text{on}}}]$, where $\mathbf{M_{\text{on}}}$ masks out the garment area in the person image $\mathbf{X_{p}}$ while preserving the garment image $\mathbf{X_{g}}$. For the try-off task, we use $\mathbf{M}=[\mathbf{1}\mid\mathbf{0}]$, masking the entire garment image and leaving the person image unmasked. The masked input is obtained as $\mathbf{X_{\text{masked}}}=\mathbf{X}\odot(1-\mathbf{M})$, where $\odot$ denotes element-wise multiplication. Our architecture follows a latent diffusion framework [4, 18], where denoising is performed in the latent space.
The full and masked images are encoded by a frozen encoder $\mathcal{E}$ into latent representations $\mathbf{z}_{0}=\mathcal{E}(\mathbf{X})$ and $\mathbf{z}_{\text{c}}=\mathcal{E}(\mathbf{X_{\text{masked}}})$. The mask $\mathbf{M}$ is downsampled via a pixel-unshuffle [70] operation to match the latent resolution, producing $\mathbf{M}_{\text{c}}$. These three components $\mathbf{z}_{0},\mathbf{z}_{\text{c}}$, and $\mathbf{M}_{\text{c}}$ are concatenated along the channel dimension, passed through a token embedding layer, and then used as input to the DiT backbone. We define the task token $\tau=[\tau_{\text{mode}}\mid\tau_{\text{category}}]$ as the concatenation of a mode token $\tau_{\text{mode}}\in\{\text{on},\text{off}\}$, which specifies the generation direction, and a category token $\tau_{\text{category}}\in\{\text{upper},\text{lower},\text{full}\}$, which encodes the garment type. The combined token $\tau$ is passed to the transformer as additional conditioning.
@@
Our framework supports dynamic input layouts by leveraging the token-based representation of vision transformers [17, 16, 60], which enables the processing of images with varying spatial resolutions and aspect ratios without requiring fixed dimensions. In our pipeline, the concatenated garment–person image is tokenized and flattened, allowing the model to accommodate inputs of arbitrary shape. Consequently, the aforementioned $H$ and $W$ need not remain fixed throughout the pipeline. To support batch training with variable-length sequences, we insert padding tokens as needed and standardize the sequence length to a fixed maximum $N_{\max}$. Additionally, we also use Rotary Position Embedding (RoPE) [74] to robustly handle inputs with diverse aspect ratios and spatial configurations. This design provides the foundation for scalable virtual try-on, enabling Voost to handle diverse poses, aspect ratios, and compositions within a single batch. By jointly modeling task direction and garment category, it supports flexible multitask learning and mitigates inductive biases introduced by the rigid preprocessing of fixed-resolution pipelines.
@@
We adopt a flow matching formulation [48], where the model learns a time-dependent velocity field that transports samples from the data distribution to noise along a continuous path. Let $\mathbf{z}_{0}$ be a data latent and $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ be a sampled noise latent. We define a trajectory $\mathbf{z}_{t}$ between them and train the denoising model $\epsilon(\cdot)$ to predict the velocity $\frac{d\mathbf{z}_{t}}{dt}$ at each intermediate point. The unified training objective is given by:
@@
In our case, we adopt the rectified flow formulation [50, 18], where the trajectory is defined as a straight line between $\mathbf{z}_{0}$ and $\mathbf{z}_{1}$:
@@ 3.3 Inference-time Refinement
Here, $d$ is the key, query, value vector dimension used in scaled dot-product attention [77]; $N_{\text{infer}}$ and $N_{\text{train}}$ are the total token counts at inference and training; $N_{\text{mask}}$ and $N_{\text{garment}}$ denote the masked and garment region tokens. The hyperparameters $\alpha$ and $\beta$ control the influence of the global and relative scaling terms, respectively, while $c$ is a fixed positive constant that stabilizes the computation when $N_{\text{mask}}$ is small. The global term stabilizes attention across token lengths, while the relative term adapts to spatial imbalance between masked and garment regions. This improves robustness to layout shifts and maintains consistent attention, particularly when the mask covers a small proportion relative to the garment (Fig. 3).
@@
Figure 4: Qualitative comparison of try-on results with existing try-on methods [41, 14, 15, 91]. Best viewed in color and under zoom.
@@ 4.1 Datasets and experimental setup
We evaluate our method on two standard benchmarks: DressCode [55] and VITON-HD [13] datasets, using both qualitative and quantitative metrics. Each dataset contains high-resolution image pairs of in-shop garments and corresponding person images. We use the official train/validation/test splits provided by both datasets. All outputs are generated at a resolution of $1024\times 768$. Our model is trained on $16\times$ NVIDIA H100 GPUs and evaluated on a single NVIDIA A100 GPU. To assess generalization, we also present qualitative results on in-the-wild images.
@@ 4.2 Qualitative comparison
Fig. 4 and Fig. 5 show qualitative comparisons between our method and state-of-the-art approaches [41, 14, 15, 91] on the VTON task. Given the increasing interest in practical applications of virtual try-on, we also include comparisons with commercial models [23, 58, 43, 57, 61] in Fig. 8.
Importantly, our unified model handles try-on and try-off with a single set of parameters, without task-specific retraining. Try-off results are shown in Fig. 6 and Fig. 10, where we compare our outputs to those of existing methods [78, 82]. Across both tasks, our method produces more coherent and photorealistic results, benefiting from shared garment-body reasoning and joint training. In particular, our model demonstrates strong robustness on in-the-wild images with diverse poses, backgrounds, and lighting conditions, consistently outperforming existing methods in these challenging scenarios. We provide additional qualitative results and comparisons in the supplementary material.

### v2

@@ 1 Introduction
Figure 1: Attention map comparison— CatVTON [chong2024catvtonconcatenationneedvirtual] shows dispersed attention unrelated to the query point, indicating weak spatial grounding. In contrast, our model produces sharply localized maps that align well with the corresponding garment regions, demonstrating stronger relational understanding.
@@
Early approaches [Han_2018_CVPR, Yang_2020_CVPR_gan, ge2021parser_gan, choi2021viton] have adopted warping-based strategies, often combined with generative modeling, but have struggled to maintain garment fidelity and structural consistency.
Effectively adapting T2I diffusion models [rombach2022high, esser2024scalingrectifiedflowtransformers] to virtual try-on requires strong image-based conditioning and accurate modeling of garment–person correspondence. To this end, various efforts have been made to improve spatial alignment and relational understanding. These include incorporating auxiliary objectives [kim2024stableviton, leffa], leveraging reference networks [Zhu_2023_CVPR_tryondiffusion, idm-vton, xu2024ootdiffusion], or injecting conditioning signals through spatial concatenation [chong2024catvtonconcatenationneedvirtual]. Despite these efforts, existing methods often fail to establish precise garment–person correspondence. As illustrated in Fig. 1, the dispersed attention patterns reflect limited relational understanding, often resulting in lower visual quality and failure to faithfully reconstruct garment details.
@@
Remarkably, our unified diffusion transformer achieves state-of-the-art performance on both VTON and VTOFF benchmarks, surpassing strong task-specific baselines [kim2024stableviton, idm-vton, leffa, chong2024catvtonconcatenationneedvirtual, tryoffdiff, tryoffanyone] optimized individually for each task. We provide extensive qualitative and quantitative analyses that validate the effectiveness of our unified approach and its ability to model garment–person interaction across diverse scenarios.
@@ 2 Related work
Diffusion models have rapidly evolved to support image-conditioned generation beyond the text-to-image setting [rombach2022high, podell2024sdxl]. Early methods such as Textual Inversion [gal2023an] and DreamBooth [ruiz2023dreambooth] enabled concept-specific synthesis via embedding injection or fine-tuning, but incurred high computational costs. Later approaches improved efficiency by leveraging pre-trained encoders like CLIP [radford2021learning] and DINO [caron2021emerging, oquab2024dinov], as seen in IP-Adapter [ye2023ip-adapter] and AnyDoor [chen2024anydoor], which use image embeddings to guide generation. To improve spatial consistency, MasaCtrl [cao2023masactrl] introduced mutual self-attention to preserve object identity. This was extended by reference-based methods [Hu_2024_CVPR, chen2024zero, xu2024ootdiffusion, idm-vton, tian2024emo] that inject structured visual features through dedicated networks. More recently, concatenation-based methods [chong2024catvtonconcatenationneedvirtual, lhhuang2024iclora, tan2024ominicontrol] have emerged as lightweight alternatives that directly feed conditioning signals without auxiliary modules. Additionally, image-conditioned diffusion has expanded beyond image-to-image generation to video [guo2023animatediff, Hu_2024_CVPR, blattmann2023stable, shi2024motion] and 3D synthesis [liu2023zero, kwak2024vivid, voleti2024sv3d], broadening its applicability.
@@
Early virtual try-on systems primarily relied on warping-based methods [wang2018toward, han2019clothflow, ge2021parser], where the main challenge was to deform and align the garment image to match the target person’s pose and body shape. With the emergence of Generative Adversarial Networks (GANs) [goodfellow2014generative, karras2019style, karras2020analyzing], a series of GAN-based approaches [Han_2018_CVPR, xie2022pastagan, Men_2020_CVPR_gan, issenhuth2020not, xie2023gpvton_gan, bai2022singlestagevirtualtryon_gan, ge2021parser_gan, Yang_2020_CVPR_gan, choi2021viton] achieved more realistic image synthesis through adversarial learning. However, these models still struggled to preserve garment details and maintain structural consistency.
With the advancement of large-scale text-to-image diffusion models [ramesh2021zero, rombach2022high, saharia2022photorealistic, podell2024sdxl, esser2024scalingrectifiedflowtransformers, flux2024], recent works [Zhu_2023_CVPR_tryondiffusion, baldrati2023multimodal, morelli2023ladi, gou2023taming_dci, idm-vton, xu2024ootdiffusion, zeng2024cat, shen2025imagdressing, chong2024catvtonconcatenationneedvirtual, leffa] have adopted pre-trained diffusion backbones as strong generative priors for virtual try-on. A key challenge remains: effectively guiding the model to learn the complex relationship between the garment and the target person. To address this, various techniques have been proposed to improve spatial correspondence and visual fidelity—such as auxiliary loss functions [kim2024stableviton, leffa], dual or reference networks [Zhu_2023_CVPR_tryondiffusion, idm-vton, chen2024wear], warping [morelli2023ladi, gou2023taming_dci, xie2023gpvton_gan, leffa], external image encoders [kim2024stableviton, idm-vton, xu2024ootdiffusion], and spatial concatenation [chong2024catvtonconcatenationneedvirtual]. However, many of these approaches rely on weak garment–person coupling or introduce additional components that may degrade generation quality.
While most existing work focuses on virtual try-on, a few studies [tryoffdiff, tryoffanyone] have explored virtual try-off, the inverse of virtual try-on, which aims to predict a clean garment image from a person wearing it. This task is inherently challenging due to occlusion, wrinkles, and deformation, and similarly requires strong relational modeling. We propose a unified DiT-based framework that jointly learns virtual try-on and try-off using a shared spatial conditioning layout. This design enables robust bidirectional garment–person modeling and remains scalable across diverse fashion imagery.
@@ 3.1 Preliminary
Diffusion models [ho2020denoising, song2021scorebased, rombach2022high] generate images by denoising a sequence of latent representations corrupted over time. Given a latent $\mathbf{z}_{t}$, the network $\epsilon_{\theta}$ is trained to recover the clean signal $\mathbf{z}_{0}$ or the noise $\boldsymbol{\epsilon}$, conditioned on external input $\mathbf{y}$. In general, the denoising step follows:
@@
where ${\Phi}(\cdot)$ denotes an integration step, such as DDIM [song2021denoising], DPM-Solver [lu2022dpm], or deterministic alternatives like flow matching [lipman2023flowmatchinggenerativemodeling]. In virtual try-on, $\mathbf{y}$ typically includes a garment image and a masked person image, optionally with pose or parsing maps [cao2017realtime, guler2018densepose]. We adopt a unified DiT [Peebles2022DiT] that handles both try-on and try-off tasks. Details follow in Sec. 3.2.
@@
A key challenge in virtual try-on is aligning the garment structure with the target body while preserving visual fidelity. We assess this correspondence by analyzing transformer attention maps at specific spatial queries. As shown in Fig.1, CatVTON [chong2024catvtonconcatenationneedvirtual] exhibits dispersed attention across unrelated areas, indicating weak spatial grounding, which can reduce structural fidelity and impair detail preservation. In contrast, attention maps produced by our model are sharply localized and semantically consistent, suggesting stronger garment–target alignment. This motivates our bidirectional training strategy, where jointly learning try-on and try-off within a single model improves spatial alignment.
@@ 3.2 Try-On & Off via Unified Transformer
Let $\mathbf{X_{g}}\in\mathbb{R}^{H\times W\times 3}$ denote the standalone garment image, and $\mathbf{X_{p}}\in\mathbb{R}^{H\times W\times 3}$ the target person image. We construct the input by horizontally concatenating the two, forming a combined image $\mathbf{X}=[\mathbf{X_{g}}\mid\mathbf{X_{p}}]\in\mathbb{R}^{H\times 2W\times 3}$. Task-specific inpainting regions are defined by a binary mask $\mathbf{M}\in\{0,1\}^{H\times 2W}$, applied in the image space prior to encoding. For the try-on task, the mask is defined as $\mathbf{M}=[\mathbf{0}\mid\mathbf{M_{\text{on}}}]$, where $\mathbf{M_{\text{on}}}$ masks out the garment area in the person image $\mathbf{X_{p}}$ while preserving the garment image $\mathbf{X_{g}}$. For the try-off task, we use $\mathbf{M}=[\mathbf{1}\mid\mathbf{0}]$, masking the entire garment image and leaving the person image unmasked. The masked input is obtained as $\mathbf{X_{\text{masked}}}=\mathbf{X}\odot(1-\mathbf{M})$, where $\odot$ denotes element-wise multiplication. Our architecture follows a latent diffusion framework [flux2024, esser2024scalingrectifiedflowtransformers], where denoising is performed in the latent space.
The full and masked images are encoded by a frozen encoder $\mathcal{E}$ into latent representations $\mathbf{z}_{0}=\mathcal{E}(\mathbf{X})$ and $\mathbf{z}_{\text{c}}=\mathcal{E}(\mathbf{X_{\text{masked}}})$. The mask $\mathbf{M}$ is downsampled via a pixel-unshuffle [shi2016real] operation to match the latent resolution, producing $\mathbf{M}_{\text{c}}$. These three components $\mathbf{z}_{0},\mathbf{z}_{\text{c}}$, and $\mathbf{M}_{\text{c}}$ are concatenated along the channel dimension, passed through a token embedding layer, and then used as input to the DiT backbone. We define the task token $\tau=[\tau_{\text{mode}}\mid\tau_{\text{category}}]$ as the concatenation of a mode token $\tau_{\text{mode}}\in\{\text{on},\text{off}\}$, which specifies the generation direction, and a category token $\tau_{\text{category}}\in\{\text{upper},\text{lower},\text{full}\}$, which encodes the garment type. The combined token $\tau$ is passed to the transformer as additional conditioning.
@@
Our framework supports dynamic input layouts by leveraging the token-based representation of vision transformers [dosovitskiy2021an_vit, dehghani2023patch, Peebles2022DiT], which enables the processing of images with varying spatial resolutions and aspect ratios without requiring fixed dimensions. In our pipeline, the concatenated garment–person image is tokenized and flattened, allowing the model to accommodate inputs of arbitrary shape. Consequently, the aforementioned $H$ and $W$ need not remain fixed throughout the pipeline. To support batch training with variable-length sequences, we insert padding tokens as needed and standardize the sequence length to a fixed maximum $N_{\max}$. Additionally, we also use Rotary Position Embedding (RoPE) [su2024roformer] to robustly handle inputs with diverse aspect ratios and spatial configurations. This design provides the foundation for scalable virtual try-on, enabling Voost to handle diverse poses, aspect ratios, and compositions within a single batch. By jointly modeling task direction and garment category, it supports flexible multitask learning and mitigates inductive biases introduced by the rigid preprocessing of fixed-resolution pipelines.
@@
We adopt a flow matching formulation [lipman2023flowmatchinggenerativemodeling], where the model learns a time-dependent velocity field that transports samples from the data distribution to noise along a continuous path. Let $\mathbf{z}_{0}$ be a data latent and $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ be a sampled noise latent. We define a trajectory $\mathbf{z}_{t}$ between them and train the denoising model $\epsilon(\cdot)$ to predict the velocity $\frac{d\mathbf{z}_{t}}{dt}$ at each intermediate point. The unified training objective is given by:
@@
In our case, we adopt the rectified flow formulation [liu2022flow, esser2024scalingrectifiedflowtransformers], where the trajectory is defined as a straight line between $\mathbf{z}_{0}$ and $\mathbf{z}_{1}$:
@@ 3.3 Inference-time Refinement
Here, $d$ is the key, query, value vector dimension used in scaled dot-product attention [vaswani2017attention]; $N_{\text{infer}}$ and $N_{\text{train}}$ are the total token counts at inference and training; $N_{\text{mask}}$ and $N_{\text{garment}}$ denote the masked and garment region tokens. The hyperparameters $\alpha$ and $\beta$ control the influence of the global and relative scaling terms, respectively, while $c$ is a fixed positive constant that stabilizes the computation when $N_{\text{mask}}$ is small. The global term stabilizes attention across token lengths, while the relative term adapts to spatial imbalance between masked and garment regions. This improves robustness to layout shifts and maintains consistent attention, particularly when the mask covers a small proportion relative to the garment (Fig. 3).
@@
Figure 4: Qualitative comparison of try-on results with existing try-on methods [kim2024stableviton, idm-vton, chong2024catvtonconcatenationneedvirtual, leffa]. Best viewed in color and under zoom.
@@ 4.1 Datasets and experimental setup
We evaluate our method on two standard benchmarks: DressCode [morelli2022dress] and VITON-HD [choi2021viton] datasets, using both qualitative and quantitative metrics. Each dataset contains high-resolution image pairs of in-shop garments and corresponding person images. We use the official train/validation/test splits provided by both datasets. All outputs are generated at a resolution of $1024\times 768$. Our model is trained on $16\times$ NVIDIA H100 GPUs and evaluated on a single NVIDIA A100 GPU. To assess generalization, we also present qualitative results on in-the-wild images.
@@ 4.2 Qualitative comparison
Fig. 4 and Fig. 5 show qualitative comparisons between our method and state-of-the-art approaches [kim2024stableviton, idm-vton, chong2024catvtonconcatenationneedvirtual, leffa] on the VTON task. Given the increasing interest in practical applications of virtual try-on, we also include comparisons with commercial models [gemini2, chatgpt4o, kolors, vella, piccopilot] in Fig. 8.
Importantly, our unified model handles try-on and try-off with a single set of parameters, without task-specific retraining. Try-off results are shown in Fig. 6 and Fig. 10, where we compare our outputs to those of existing methods [tryoffdiff, tryoffanyone]. Across both tasks, our method produces more coherent and photorealistic results, benefiting from shared garment-body reasoning and joint training.
In particular, our model demonstrates strong robustness on in-the-wild images with diverse poses, backgrounds, and lighting conditions, consistently outperforming existing methods in these challenging scenarios. We provide additional qualitative results and comparisons in the supplementary material.